the hypotenuse of this triangle is 17 meters. if there are no decimal answers, what would the actual measures of sides a and b have to be?
Found 3 solutions by Alan3354, ewatrrr, richard1234:
Answer by Alan3354(69443)

8 & 15 meters

Answer by ewatrrr(24785)


Hi,

hypotenuse of this triangle is 17 meters...
what would the actual measures of sides a and b have to be if a and b were whole numbers
using the Pythagorean Theorem
17m^2 = a^2 + b^2
IF a = 15m AND b = 8m
17m^2 = 289m = 225m + 64m

Answer by richard1234(7193)

8 meters and 15 meters, if the sides are integers. You'll have to memorize a few Pythagorean triples to get this one quickly.